OpenStudy (anonymous):

Determine which of the following statements is true concerning the values described in column #1 and column #2. Column #1 The x-coordinate of the vertex of the equation y = 2x2 − 4x + 12 Column #2 The x-coordinate of the vertex of the equation y = 4x2 + 8x + 3

OpenStudy (anonymous):

Possible Answers Possible Answers 1) The value found in column #1 is greater than the value found in column #2. 2)The value found in column #1 is less than the value found in column #2. 3) The value found in column #1 is equivalent to the value found in column #2 4) The relationship between column #1 and column #2 cannot be determined by the given info

OpenStudy (anonymous):

before we answer the questions, why don't we find the coordinates of the vertex for each? i

OpenStudy (anonymous):

the first coordinate of the vertex of \(ax^2+bx+c\) is \(-\frac{b}{2a}\) compute the two numbers

OpenStudy (anonymous):

for example, in \(y = 2x^2 − 4x + 12\) we have \(a=2,b=-4\) and so \[-\frac{b}{2a}=-\frac{-4}{2\times 2}=-1\]

OpenStudy (anonymous):

sorry that was wrong it is \[-\frac{2}{2a}=-\frac{-4}{2\times 2}=1\] not \(-1\) that was my mistake

OpenStudy (anonymous):

Ok I think i'm following you so far.

OpenStudy (anonymous):

good then you only have one more to do compute the vertex the same way for \[y = 4x^2 + 8x + 3 \] that is, compute \(-\frac{b}{2a}\) for \(a=4,b=8\) let me know what you get

OpenStudy (anonymous):

and this would be where I lost you lol

OpenStudy (anonymous):

I believe the answer is the first choice. Is this right or wrong?

OpenStudy (anonymous):

did you calculate \[-\frac{8}{2\times 4}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

i hope you got \(-1\) so you are right, the first answer is correct,

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Ok thank you!